OpenStudy (anonymous):

Factor: 9x^2-5x+3=0 3x^2+4x-2=0 x^2-6x+10=0 x^2+9x+9=0 And I need help with this question: If a baseball player hits a baseball from 4 feet off the ground with an initial velocity of 64 feet per second, how long will it take the baseball to hit the ground? Use the equation h = -16t2 + 64t + 4. Fan/medal for help. :)

OpenStudy (anonymous):

That's not one of my answer choices.

OpenStudy (anonymous):

2 plus or minus square root of 17 end root over 2 quantity of 2 plus or minus square root of 17 all over 2 2 plus or minus 4 square root of 17 quantity of 16 plus or minus square root of 17 all over 2 Probably should have put those there. ^^

OpenStudy (phi):

how long will it take the baseball to hit the ground? when it hits the ground, it will have a height of 0. solve for t \[-16t^2 + 64t + 4 = 0 \]

OpenStudy (anonymous):

And then you solve the quadratic equation that is: -16t^2+64t+4=0

OpenStudy (phi):

yes, use the quadratic formula

OpenStudy (anonymous):

How do you factor it?

OpenStudy (phi):

first, notice we can divide all terms by 4, to make the numbers smaller \[ -4 t^2 + 16t+1= 0 \]

OpenStudy (anonymous):

Would it look like: -4(4x^2+16+1)=0?

OpenStudy (anonymous):

64/4=16?

OpenStudy (phi):

almost, \[ -4( 4t^2 -16t -1) = 0 \\ 4t^2 -16t -1 = 0 \]

OpenStudy (phi):

do you know the quadratic formula? or how to complete the square ?

OpenStudy (anonymous):